OpenStudy (anonymous):

solve √4x+2 +1= √5x+4 please help :)

OpenStudy (marigirl):

can u clarify the question plz \[\sqrt{4x+2}+1=\sqrt{5x+4}\]

OpenStudy (anonymous):

nevermind, i solved it. thank you anyway!

OpenStudy (marigirl):

Well done!

OpenStudy (anonymous):

thank you so much!!